# Direct Quantized Training of Language Models with Stochastic Rounding

### v1

@@ Abstract
Although recent quantized Large Language Models (LLMs), such as BitNet, have paved the way for significant reduction in memory usage during deployment with binary or ternary weights, training these models still demands substantial memory footprints. This is partly because high-precision (i.e., unquantized) weight matrices required for straight-through estimation must be maintained throughout the whole training process. To address this, we explore the potential of directly updating the quantized low-precision weight matrices without relying on the straight-through estimator during backpropagation, thereby saving memory usage during training. Specifically, we employ a stochastic rounding technique to minimize information loss caused by the use of low-bit weights throughout training. Experimental results on our LLaMA-structured models indicate that (1) training with only low-precision weights is feasible even when they are constrained to ternary values, (2) extending the bit width to 8 bits results in only a 5% loss degradation compared to BitNet b1.58 while offering the potential for reduced memory usage during training, and (3) our models can also perform inference using ternary weights, showcasing their flexibility in deployment. [^1] Work in progress. Code is available at https://github.com/KYuuto1006/DQT
@@ 1 Introduction
Large Language Models (LLMs) have become a promising solution for a wide range of Natural Language Processing (NLP) tasks including machine translation (Xu et al., 2024a; Wu et al., 2024), summarization (Zhang et al., 2024a), reasoning (Wei et al., 2024; Kojima et al., 2024; Miao et al., 2024) and question answering (OpenAI, 2022, 2023). However, their development is challenged by the need for vast datasets and substantial computational resources, especially when the size of current LLMs continually grows larger (Duan et al., 2024).
@@
Quantization, which involves converting high-precision parameter matrices into lower-precision formats, emerges as an effective approach for enabling resource-efficient LLMs (Egashira et al., 2024). Traditional quantization methods can be divided into two categories: Post Training Quantization (PTQ) and Quantization Aware Training (QAT). PTQ reduces the bit precision of weight matrices in an already pre-trained LLM (Banner et al., 2019; Frantar et al., 2023), while QAT incorporates quantization during training, enabling the model to adapt to low-bit precision throughout the learning process (Jacob et al., 2018). Numerous studies have shown that models quantized through QAT generally achieve better performance than those using PTQ because PTQ may lead to a mismatch between the high-precision model’s learned representations and the constraints of low-bit precision (Chen et al., 2024; Liu et al., 2024).
@@
Figure 1: Comparison of the training process for BitNet and our modified one. Upper: The training process for BitNet, where the original high precision weights are updated with the straight-through estimator in backward process. Lower: We directly update the low precision weights with stochastic rounding, eliminating the need to quantize the weight matrices in each training step and keeping weight matrices always ternary.
@@
Recently, a QAT method, BitNet (Wang et al., 2023; Ma et al., 2024), makes the first attempt to quantize full-precision (FP32) transformers into binary or ternary models from scratch, with weight matrices constrained to values of {-1, 0, 1} while maintaining competitive performances with unquantized ones. We illustrate the training process of traditional QAT methods such as BitNet in Figure 1 (a). After computing the loss based on the quantized weights and inputs, the loss is backpropagated to update the original high-precision weights via the straight-through estimator (STE) (Bengio et al., 2013). These weights are then re-quantized in each training step. This iterative process is necessary because quantization itself is not differentiable (Chen et al., 2019), requiring special gradient accumulation methods. As a result, the high-precision weight matrices are always maintained throughout the whole training process, which makes traditional QAT inefficient and takes up a lot of extra memory footprints.
Particularly, when scaling to LLMs, this problem may pose significant challenges to QAT’s practical deployment, as maintaining larger weight matrices in LLMs becomes more resource-intensive. For example, maintaining the weights of a 1B LLM alone would require 4GB of memory in the FP32 format, whereas maintaining ternary weights reduces this to 0.2GB. This requirement for memory footprints limits the accessibility of QAT techniques for researchers and organizations with limited computational resources.
@@
In order to address these challenges, we explore Direct Quantized Training (DQT), a modified QAT approach for language models that maintains only low-precision weights throughout the training process. DQT eliminates the reliance on STE by directly updating the quantized low-precision weights during backpropagation, as shown in Figure 1 (b). Specifically, we use stochastic rounding (Von Neumann and Goldstine, 1947), a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values, thereby preserving the low-precision format of the weight matrices after backpropagation and minimizing the information loss caused by using only low-bit weights. The DQT method keeps all weight matrices fixed at n-bit precision throughout the entire training process. With DQT, the model no longer needs to quantize all high-precision weight matrices at each training step, theoretically accelerating the QAT process and reducing memory usage. More importantly, traditional QAT often introduces extra memory overhead due to the updates on high-precision weights, which limits its practicability in training general LLMs. In contrast, the light-weighted memory dependency of DQT enables quantization’s application even for scenarios where computational resources are constrained.
@@
Experimental results on our LLaMA-structured models trained with DQT demonstrate that (1) models trained with DQT are capable of converging even when weight matrices are constrained to ternary values; (2) with 8-bit DQT, our models can achieve performance levels competitive with BitNet b1.58, showing the feasibility of DQT; and (3) it is also possible for models trained with DQT to utilize only ternary weights during inference while delivering performance comparable to BitNet. We assume that DQT could provide new insights on addressing the computational challenges posed by traditional QAT.
@@ 2 Related Works
Quantization for deep neural networks has a history spanning nearly ten years, with researchers initially compressing networks to reduce memory usage and computational load while maintaining accuracy (Rastegari et al., 2016; Hubara et al., 2018). Recent quantization methods for LLMs can be divided into two categories: Post Training Quantization (PTQ) and Quantization Aware Training (QAT).
2.1 Post Training Quantization
@@
2.2 Quantization Aware Training
To bridge the gap between high-precision parameter training and quantization, QAT incorporates the quantization of model parameters during the training process. However, the substantial computational cost of applying QAT to LLMs has hindered its widespread exploration.
The first application of QAT to LLMs comes from Liu et al. (2024), who propose a data-free distillation-based method and quantize the model to 4 bits. Xu et al. (2024b) further expand distillation-based methods to binary quantization by introducing limited trainable vectors. More recently, BitNet (Wang et al., 2023; Ma et al., 2024) is proposed, achieving training from scratch QAT with weight values constrained to {-1,0,1}.
However, due to the non-differentiability of the quantization process, special gradient approximation methods like the straight-through estimator (STE) (Bengio et al., 2013) are commonly employed during training. While effective, this approach often results in slower training processes and increased computational memory overhead since high-precision weights are always maintained during the training process, limiting its scalability to larger models.
These inefficiencies become particularly pronounced when scaling QAT to larger models, such as those in the billion-parameter range, where the memory and computational demands can quickly surpass the capabilities of standard GPUs, limiting the practical application of QAT for real-world use cases. To address this, in this work we explore the potential for more efficient and practical QAT methods that significantly reduce memory usage compared to traditional QAT approaches.
@@ 3.1 Stochastic Rounding
The idea of stochastic rounding (SR) originates from Von Neumann and Goldstine (1947), and is initially used for reducing bias in numerical computations and recently for deep learning models (Gupta et al., 2015). It is a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values. Given a high precision value $x$, stochastic rounding can be defined as the following equation Gupta et al. (2015); Zhang et al. (2024b):
@@ 3.2 Modified Training Process
where $\mathrm{Clip}()$ function assures that all the values are in the range of $[Q_{n},Q_{p}]$ and $\mathrm{Round}()$ returns the nearest integer. This allows us to constrain the quantized $\tilde{W}$ to $n$-bits. While for the inputs and activations, we follow the settings introduced in BitNet (Wang et al., 2023; Ma et al., 2024) and quantize them into 8 bits.
@@
to make sure it maintains $n$-bits without requiring the retention of high-precision weights, thus getting rid of the straight-through estimator and skipping the process from Equation (2) to Equation (4). During our modified DQT, we directly replace $\tilde{W}$ with $\widetilde{W}$ and proceed to the next training step, ensuring that the weight matrices are always constrained to $n$-bits throughout training. This makes the biggest difference from traditional QAT methods.
@@ 4 Experiments
In this section, we begin by detailing our training dataset and implementation setup. Next, we compare models trained with DQT to FP32 models in Section 4.4. We move on to discuss the impact of altering the bit width involved in DQT in Section 4.5. Finally, in Section 4.6, we conduct a fair comparison between our DQT models and BitNet b1.58, both utilizing ternary weight matrices.
4.1 Training Dataset
We use the English Wikipedia dataset (20231101.en) [^2] https://huggingface.co/datasets/wikimedia/wikipedia to pre-train baselines and our models. The maximum length of training data is set to 512. For texts longer than 512, we split them into individual texts while for those shorter than 512, padding is applied. As a result, we acquire a dataset with 14 million sentences from the original 6.4 million examples.
4.2 Baselines
@@
We choose two types of baselines for comparing with our DQT. The first is a reproduced full-precision (FP32) model. We compare DQT to it to demonstrate the effectiveness of our approach with reduced bits. The second baseline is the reproduced BitNet b1.58, a well-known QAT method that employs ternary weights. To ensure a fair comparison with BitNet, we constrain our DQT models to use only ternary weights during inference.
4.3 Implementation Details
We train all our models using fake (simulated) quantization on NVIDIA A100 GPUs to remain hardware-agnostic, without leveraging true low-precision computation. We follow the structure of a 12-layer LLaMA Touvron et al. (2023); Dubey et al. (2024) and initiate our model with the hidden size and the intermediate size set to 768, 2048, respectively. The number of attention heads is set to 12 and the maximum position embeddings follow the maximum input length, which is set to 512. Thus, the total parameters for our model and baselines equal to around 130 million. For the baselines, AdamW is utilized as the optimizer while for DQT models, modified AdamW with stochastic rounding is employed. The cosine scheduler is applied with a 2000 step warm-up. We pre-train all models for one epoch under our dataset on 4 A100 80G GPUs. The batch size is set to 64 on each GPU with no gradient accumulation. The learning rate is grid-searched from {1e-5, 1e-4, 5e-4, 1e-3} for each model. As for the tokenizer, we directly utilize a publicly released pre-trained one  [^3] https://huggingface.co/1bitLLM/bitnet_b1_58-large and do not update it during the training process. We train all of our models under FP32 with mixed precision enabled.
4.4 Training Comparison
Figure 2: Comparison of our DQT and other baselines. The horizontal axis represents the training steps. Optimal learning rates are individually selected for each model, specifically 1e-4 for BitNet and the FP32 model, 1e-3 for DQT 1.58-bit, and 5e-4 for DQT 8-bit implementations.
We present the training loss for DQT variants and our reproduced LLaMA, BitNet b1.58 in Figure 2, using the optimal learning rate for each model. In Figure 2, the blue line indicates the ternary implementation of DQT, demonstrating that it is capable of convergence.
When comparing our DQT ternary variant to the reproduced FP32 model, a performance gap is observed. This is expected, as the DQT model operates with only 1.58-bit information during training, limiting its capacity compared to the FP32 model. However, when longer bit-width models are employed, quantizing weight parameters into 8 bits and maintaining only these 8-bit values during training, the performance becomes much more comparable to the FP32 model.
As for the comparison with BitNet b1.58, since our DQT ternary variant does not rely on high-precision weights during training, it results in lower performance. However, as only ternary weights are maintained throughout the process, this approach potentially requires less memory than BitNet. Even when 8-bit quantization is enabled in DQT, the high-precision information required for BitNet exceeds that of our DQT approach.

### v2

@@ Abstract
Although recent quantized Large Language Models (LLMs), such as BitNet, have paved the way for significant reduction in memory usage during deployment with binary or ternary weights, training these models still demands substantial memory footprints. This is partly because high-precision (i.e., unquantized) weights required for straight-through estimation must be maintained throughout the whole training process. To address this, we explore directly updating the quantized low-precision weights without relying on straight-through estimation during backpropagation, aiming to save memory usage during training. Specifically, we employ a stochastic rounding technique to minimize the information loss caused by the use of low-bit weights throughout training. Experimental results on our LLaMA-structured models of various sizes indicate that (1) training with only low-precision weights is feasible even when they are constrained to ternary values; (2) extending the bit width to 8 bits achieves performance on par with BitNet b1.58; (3) our models remain robust to precision scaling and memory reduction, showing minimal performance degradation when moving from FP32 to lower-memory environments (BF16/FP8); and (4) our models also support inference using ternary weights, showcasing their flexibility in deployment.  [^1] Work in progress. Code is available at https://github.com/KYuuto1006/DQT
@@ 1 Introduction
Large Language Models (LLMs) have become a promising solution for a wide range of Natural Language Processing (NLP) tasks, including machine translation (Xu et al., 2024a; Wu et al., 2024), summarization (Zhang et al., 2024a), reasoning (Wei et al., 2024; Kojima et al., 2024; Miao et al., 2024) and question answering (OpenAI, 2022, 2023). However, their development is challenged by the need for vast datasets and substantial computational resources, especially as the size of current LLMs continually grows larger (Duan et al., 2024).
@@
Quantization, which involves converting high-precision parameter matrices into lower-precision formats, has emerged as an effective approach for enabling resource-efficient LLMs (Egashira et al., 2024). Traditional quantization methods can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ reduces the bit precision of weight matrices in an already pre-trained LLM (Banner et al., 2019; Frantar et al., 2023), while QAT incorporates quantization during training, enabling the model to adapt to low-bit precision throughout the learning process (Jacob et al., 2018).
@@
Figure 1: Comparison of the training process for BitNet and our modified one. Upper: The training process for BitNet, where the original high precision weights are updated with the straight-through estimator in backward process. Lower: We directly update the low precision weights with stochastic rounding, eliminating the need to quantize the weight matrices in each training step and keeping weight matrices always at low-bit. For clarity, we illustrate our method using ternary examples. We present an example of 8-bit quantization in Appendix, Figure 8.
@@
Recently, a QAT method, BitNet (Wang et al., 2023; Ma et al., 2024), has shown the feasibility of quantizing full-precision (FP32) transformers into binary or ternary models from scratch, while maintaining competitive performance with unquantized ones. We illustrate the training process of traditional QAT methods such as BitNet in Figure 1 (a). After computing the loss based on the quantized weights and inputs, the loss is backpropagated to update the original high-precision weights via the straight-through estimator (STE) (Bengio et al., 2013). These weights are then re-quantized in each training step. This iterative process is necessary because quantization itself is not differentiable (Chen et al., 2019), requiring special gradient accumulation methods. As a result, the high-precision weight matrices are always maintained throughout the whole training process, which makes traditional QAT inefficient and takes up a lot of extra memory footprints. For example, maintaining the weights of a 1B LLM alone would require 4GB of memory in the FP32 format, whereas maintaining ternary weights reduces this to 0.2GB. This requirement for memory footprints limits the accessibility of QAT techniques for researchers and organizations with limited computational resources.
@@
In order to address these challenges, we explore Direct Quantized Training (DQT), a modified QAT approach for language models that maintains only low-precision weights throughout the training process. DQT eliminates the reliance on STE by directly updating the quantized low-precision weights during backpropagation, as shown in Figure 1 (b). Specifically, we use stochastic rounding (Von Neumann and Goldstine, 1947) to preserve the low-precision format of the weight matrices after backpropagation and minimize the information loss caused by using only low-bit weights. The DQT method keeps all weight matrices fixed at n-bit precision (INTn) throughout the entire training process and eliminates the needs to quantize high-precision weight matrices at each training step. More importantly, traditional QAT often introduces extra memory overhead due to the updates on high-precision weights, which limits its practicability in training general LLMs. In contrast, the light-weighted memory dependency of DQT enables quantization’s application even for scenarios where computational resources are constrained.
@@
Experimental results on our LLaMA-structured models ranging from 130M, 320M and 1B parameters demonstrate that (1) DQT enables model convergence even when weight matrices are constrained to ternary values; (2) with 8-bit DQT, our models can achieve performance levels competitive with BitNet b1.58, showing the feasibility of the approach; (3) DQT models exhibit robust performance against GPU memory reductions, showing their light-weighted memory dependency; and (4) inference using only ternary weights in DQT remains effective, delivering performance comparable to BitNet. In addition, we conduct an in-depth analysis of stochastic rounding and find that it helps preserve critical update signals and contributes to training stability. We assume that DQT could provide new insights on addressing the computational challenges posed by traditional QAT.
@@ 2 Related Works
Quantization for deep neural networks has a history spanning nearly ten years, with researchers initially compressing networks to reduce memory usage and computational load while maintaining accuracy (Rastegari et al., 2016; Hubara et al., 2018). Recent quantization methods for LLMs can be divided into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
2.1 Post-Training Quantization
@@
2.2 Quantization-Aware Training
To bridge the gap between high-precision parameter training and quantization, QAT incorporates the quantization of model parameters during the training process. One of the initial applications of QAT to LLMs comes from Liu et al. (2024), who propose a data-free distillation-based method and quantize the model to 4 bits. Xu et al. (2024b) further expand distillation-based methods to binary quantization by introducing limited trainable vectors. More recently, BitNet (Wang et al., 2023; Ma et al., 2024) is proposed, achieving training from scratch QAT with weight values constrained to {-1,0,1}. However, due to the non-differentiability of the quantization process, special gradient approximation methods like the straight-through estimator (STE) (Bengio et al., 2013) are commonly employed during training. While effective, this approach often results in slower training processes and increased computational memory overhead since high-precision weights are always maintained during the training process. These inefficiencies become particularly pronounced when scaling QAT to larger models, limiting the practical training of QAT in real-world use cases. To address this, in this work, we explore the potential for more efficient and practical QAT methods that have significantly lower memory dependency compared to traditional QAT approaches.
@@ 3.1 Stochastic Rounding
The idea of stochastic rounding (SR) originates from Von Neumann and Goldstine (1947), and is initially used for reducing bias in numerical computations and recently for deep learning models (Gupta et al., 2015). It is a rounding technique that probabilistically rounds values to the nearest representable precision based on their distance from those values. Given a high precision value $x$, stochastic rounding can be defined as the following equation Gupta et al. (2015); Markov et al. (2023); Zhang et al. (2024b):
@@ 3.2 Modified Training Process
where $\mathrm{Clip}()$ function assures that all the values are in the range of $[Q_{n},Q_{p}]$ and $\mathrm{Round}()$ returns the nearest integer. This allows us to constrain the quantized $\tilde{W}$ to $n$-bits (INTn). While for the inputs and activations, we follow the settings introduced in BitNet (Wang et al., 2023; Ma et al., 2024) and quantize them into 8 bits.
@@
to make sure it maintains $n$-bits (INTn) without requiring the retention of high-precision weights, thus getting rid of the straight-through estimator and skipping the process from Equation (2) to Equation (4). During our modified DQT, we directly replace $\tilde{W}$ with $\widetilde{W}$ and proceed to the next training step, ensuring that the weight matrices are always constrained to $n$-bits throughout training, which makes the biggest difference from traditional QAT.
@@ 4 Experiments
In this section, we begin by detailing our training dataset and implementation setup in Section 4.1. Section 4.2 presents a comparison between the training behavior of DQT models and other baselines. To demonstrate the robustness of DQT under reduced GPU memory environments, we evaluate its performance in FP32, BF16, and FP8 formats, as well as with memory-efficient optimizers, in Section 4.3. We then analyze the effect of varying bit widths in DQT in Section 4.4. Finally, Section 4.5 reports evaluation results on various tasks.
4.1 Implementation Setup
Model architecture.
We follow the architecture of LLaMA2 Touvron et al. (2023); Dubey et al. (2024) and initialize models with sizes ranging from 130M to 320M and 1B parameters. Detailed model specifications and configurations are provided in Appendix A.1. For the main experiments, we use AdamW Loshchilov and Hutter (2019) as the optimizer. For DQT models, we apply modified versions of the optimizers that incorporate stochastic rounding.
Baselines.
@@ 4.1 Implementation Setup
We choose two types of baselines for comparison with our DQT. The first is a reproduced full-precision (FP32) model. We compare DQT to it to demonstrate the effectiveness of our approach with reduced bits. The second baseline is the reproduced BitNet b1.58, a well-known QAT method that employs ternary weights at inference but relies on high-precision weights during training.
Datasets
We use two types of datasets to pretrain the models. The first is the English Wikipedia dataset (20231101.en) [^2] https://huggingface.co/datasets/wikimedia/wikipedia. For the 1B models, we further pre-train them using a larger dataset containing 10B tokens from the FineWeb dataset (Penedo et al., 2024) [^3] https://huggingface.co/datasets/HuggingFaceFW/fineweb/viewer/sample-10BT. We split 1% of the data as the corresponding development set. Models are trained for one epoch with a cosine scheduler applied and a 2000 step warm-up. Refer to Appendix A.1 for details.
Evaluation.
Besides reporting the training loss and perplexity results, we perform zero-shot evaluation on general language modeling tasks using the lm_eval benchmark Gao et al. (2024). Specifically, we evaluate on WinoGrande Sakaguchi et al. (2019), ARC Clark et al. (2018), PIQA Bisk et al. (2020) and SciQ Welbl et al. (2017), which span a range of reasoning and question answering tasks across diverse domains.
Figure 2: Comparison of our DQT and other baselines accross different model sizes and training datasets. The horizontal axis represents the training steps while the vertical axis represents the training loss. As model size increases, the performance of our DQT models, especially DQT 8 bit, become more comparable to and even better than the reproduced BitNet b1.58.
Hardware.
We pre-train the 130M models on 4 NVIDIA A100 80GB GPUs, the 330M models on 8 NVIDIA GH200 Grace Hopper Superchips, and the 1B models on 16 GH200 Superchips. To remain hardware-agnostic, we simulate our quantization approach under FP32, BF16, and FP8 environments. We use MS-AMP [^4] https://github.com/Azure/MS-AMP for FP8 experiments with optimization level set to MS-AMP O2. While we recognize that FP8 is still not true ternary precision, it offers a practical compromise under current hardware constraints, as done in many prior works. Refer to Appendix A.1 for more details.
4.2 Main Results
We first present the training loss for DQT variants, our reproduced LLaMA (FP32) and BitNet b1.58 in Figure 2, including different model sizes and training datasets under FP32. In all subfigures of Figure 2, the blue line represents the ternary implementation of DQT, where weight matrices are always constrained to {-1, 0, 1}. The green line denotes the 8 bit (INT8) implementation of DQT. The orange line corresponds to the reproduced BitNet b1.58, which utilizes high-precision information during training. Finally, the red line indicates the standard FP32 LLaMA implementation.
We first examine the blue lines: although there remains a performance gap compared to higher-precision models, the ternary DQT implementation still demonstrates the ability to converge. Across all model sizes and training datasets, we observe that standard FP32 models (red lines) consistently achieve the best performance, as they are not subject to any quantization. BitNet (orange lines) consistently delivers performance close to FP32, benefiting from high-precision updates during training. Finally, for our DQT 8 bit implementation (green lines), the performance gap with BitNet b1.58 narrows as model size increases. Particularly, in the 1B models, DQT 8 bit models surpass BitNet b1.58, and this trend persists even when training on the larger FineWeb dataset. A clearer non-logarithmic comparison between the 1B DQT 8-bit model and BitNet b1.58 is provided in Appendix, Figure 10.
Generally speaking, as model size increases, all models show improved performance, reflecting the general benefits of scaling. However, the performance gains of our DQT models are especially pronounced. This suggests that DQT benefits more from scaling than other quantized approaches, narrowing the gap with other baselines and, in some cases, even surpassing them. Notably, since our DQT models do not rely on high-precision weights during training, even when 8-bit quantization is used, their memory requirements are significantly lower than those of BitNet, which still depends on high-precision updates. In the next section, we demonstrate DQT’s low GPU memory dependency in detail.
4.3 Low Memory Experiments
We additionally conduct two types of experiments to evaluate DQT’s performance given reduced GPU memory, simulating resource-limited environments. First, we assess performance in low-precision formats using BF16 and FP8. Second, we examine the effect of memory-efficient optimizers, as standard AdamW maintains two high-precision states (momentum and variance) for each parameter, contributing significantly to memory usage. To demonstrate DQT’s low reliance on high-precision information, we specifically choose Adafactor (Shazeer and Stern, 2018), which eliminates the need for fully storing these additional states. Note that we conduct experiments with Adafactor in BF16 and FP8 formats, as its benefits are most pronounced in low-precision environments.
Figure 3 presents the GPU memory usage and corresponding development set loss at the end of training for 130M and 1B models trained on the Wikipedia dataset. The y-axis indicates the percentage of GPU memory used on a single GH200 GPU under our experimental settings. We explain how we acquire the percentage in Appendix A.3. As shown in the figure, BitNet models experience clear performance degradation as GPU memory usage decreases in BF16 and FP8 settings, for both 130M and 1B scales. This is expected, since BitNet still relies on high-precision information for effective weight updates. In contrast, our DQT 8-bit models maintain robust performance under reduced precision. For both 130M and 1B models, the performance drop is less than 0.1 in loss, demonstrating DQT’s resilience to lower GPU memory usage and its minimal dependence on high-precision information. This robustness persists even when applying memory-efficient optimizers such as Adafactor in both BF16 and FP8 settings. These results suggest that DQT models can be effectively trained even in memory-constrained environments, showing consistent performance across both low-precision formats and lightweight optimization methods. We believe that DQT can be further extended to even lower-precision settings while maintaining its effectiveness.
Figure 3: GPU memory usage versus loss on development set. While BitNet suffers significant performance degradation in low-precision formats, DQT demonstrates strong robustness with minimal loss increase.